# Revealing the Geometrical and Vibrational Properties of the Defects Driving the Boson Peak

### v1

@@ Abstract
Figure 4: Quadrupolar nature of defects. a Pearson correlation coefficient between coarse-grained cage-relative displacement and quadrupolar charge. b Quadrupoles associated with the cluster superimposed on a heat map of the magnitude of the coarse-grained cage-relative displacement, normalised by its average value. The figure refers to $x_{c}=1.7$ and depicts 1/12th of the system’s area. c Density map of the two largest (in modulus) eigenvalues of the deviatoric strain associated with the defects, normalised to its maximum value. The density is higher close to the $\Upsilon_{2}=-\Upsilon_{1}$ line, implying that the quadrupoles correspond to pure shear deformations. d We represent the quadrupolar deformations by attaching to each defect’s centre of mass the displacements along the principal axis ±$\pm\Upsilon_{1}{\bm{u}}_{1}$ (blue), ±$\pm\Upsilon_{2}{\bm{u}}_{2}$ (green) and ±$\pm\Upsilon_{3}{\bm{u}}_{3}$ (red). Since quadrupoles mostly correspond to pure-shear deformations, ${\bf u}_{3}$ is barely visible.
The localised defects regulate the boson peak
We test more rigorously whether quadrupolar defects regulate the boson peak by evaluating if their density $\rho_{d}$ (number per unit volume) accounts for the excess density of vibrational modes defined as $\Delta D(\omega_{\rm bp})=D(\omega_{\rm bp})-D_{\rm Debye}(\omega_{\rm bp})$, where $D_{\rm Debye}(\omega_{\rm bp})=A_{D}\omega^{D-1}$ represents Debye’s prediction, as expected if the phonon-defect interaction is weak. In making this test, it is important to note that $\rho_{d}$ is not uniquely determined, as it depends on the definition of defects used. We define defects as clusters of particles with cage-relative displacements in the top 5%. This threshold is not critical, as the SI shows that varying it does not significantly affect the number of defects or $\rho_{d}$, as it primarily impacts the defect size. Instead, the density of defects is more influenced by the minimum number of particles, $n_{\rm min}$, required for a cluster to be considered a defect. This threshold is introduced to ensure that volume-preserving deformations involve more than just a few particles.
Fig. 5(a) explores the dependence of the density of defects on $x_{c}$ in two-dimensional systems. An $n_{\rm min}$ increases, $\rho_{d}$ quickly looses any $x_{c}$ dependence. Similarly, we show in panel (b) that the excess of modes $\Delta D(\omega_{\rm bp})$ is also constant in two dimensions, showing therefore a clear correlation with the also constant $\rho_{d}$. In three dimensions, Fig. 5(c) demonstrates that the density of defects increases as $x_{c}$ decreases, consistent with the loss of stability. The excess of modes follows the same trend with $x_{c}$, with Fig. 5(d) demonstrating their proportionality. Importantly, in 2D, $\Delta D(\omega_{\rm bp})$ and $\rho_{d}$ have comparable values, besides being both $x_{c}$ independent. In 3D, we find that $\Delta D(\omega_{\rm bp})=c\rho_{d}$. The constant of proportionality $c$ depends on the specific definition of defects, but it is consistently of order 1. These findings strongly suggests that the excess of vibrational modes at the boson peak frequency originates from the quadrupolar defects we have identified.
Figure 5: Density of vibrational defects and excess density of modes. We evaluated the excess density of vibrational modes at the boson peak frequency, $\Delta D(\omega_{\rm bp})=D(\omega_{\rm bp})-D_{\rm Debye}(\omega_{\rm bp})$, $D_{\rm Debye}(\omega)$ being Debye’s density of state, and the density $\rho_{d}$ of vibrational defects containing more than $n$ particles. In two dimensions, a $\rho_{d}$ for clusters containing a minimum number of particle $n_{\rm min}$ is $x_{c}$ independent. Similarly, b, $\Delta D(\omega_{\rm bp})$ is $x_{c}$ independent. In three dimensions, c $\Delta D(\omega_{\rm bp})$ grows with $x_{c}$, and d $\rho_{d}\propto\Delta D(\omega_{\rm bp})$.
Discussion and relation to stringlets
We have identified localised defects hybridizing with plane waves at the boson peak frequency through a filtering technique based on the physical insight that localized vibrations are generally less affine than extended ones. This technique allowed us to isolate the vibrational defects and study their geometrical and vibrational properties. The vibrational defects have some degree of spatial anisotropy but are compact rather than planar or string-like. Their oscillatory motion is quadrupolar. Specifically, the defects undergo pure shear oscillatory deformations, extending along one axis while contracting along a perpendicular one in both two and three spatial dimensions. Analysis of various two- and three-dimensional systems shows that the number density of these defects quantitatively accounts for the excess of modes at the boson peak, confirming their physical relevance.
At this point, one might then wonder how our findings are compatible with the recent claims by Hu and Tanaka [25], that on the contrary identify the microscopic origin of the BP in one-dimensional string-like dynamical objects. These ‘stringlets‘ are defined as clusters of particles with a large vibrational amplitude, or equivalently, a high ‘vibrability’ $\Psi_{i}=\sum_{k}\frac{1}{\omega_{k}^{2}}\left|{\bf e}_{k,i}\right|^{2}$, where the sum extends over all modes with frequencies $\omega_{k}$ close to the boson peak frequency $|\omega_{\text{bp}}-\omega_{k}|<\Delta\omega$. We have performed the same stringlet analysis in our simulation model. As shown in Figure 6(a) by the average coordination number of the particles participating in the stringlets, we find that the clusters defined as in [25] are compact and extended objects, rather than one-dimensional, very similar to the localised defects found in our work (Figure 3(b)). In Figure 6(b), we observe a high spatial correlation between particles with vibrability in the top 5% (red) alongside those with cage-relative displacements in the top 5% (blue), indicating a high correlation between vibrational motion and cage-relative displacement at the BP frequency. These correlations are further confirmed by measuring the fraction $f$ of particles with high vibrability and cage-relative displacement, which would be $f_{\text{rand}}=0.05$ in the absence of correlations. Figure 6(c) shows that $f$ is independent of $x_{c}$ and significantly exceeds the random expectation.
Our analysis provides strong evidence that quadrupolar localised defects and high-vibrability clusters (previously identified as stringlets) are two sides of the same coin. This reconciles two of the most successful theories for the origin of the BP in glasses and reveal the ultimate nature of the defects responsible for it.
Along the lines of [15], it would be interesting to understand whether the same defective structures could be also rationalized by means of shear modulus fluctuations and whether our findings are compatible with another successful theory of the BP – heterogeneous elasticity theory [45]. Moreover, it would be fruitful to investigate the relation between the defects discussed in our manuscript and (I) the geometric charges responsible for plastic screening and anomalous elastic response [46], (II) the topological vortex-like defects recently connected to plastic soft spots [47, 48].
@@
Figure 6: Reconciling stringlets and quadrupolar defects. a The coordination number of vibrability clusters increases with $n$, following a trend similar to that of clusters formed based on cage-relative displacement (Fig. 3c). b Particles exhibiting high cage-relative displacement at the boson peak (red) and high vibrability (blue), both in the top 5%, for $x_{c}=1.7$. Strong correlations between the two are observed. c The fraction of particles displaying both large cage-relative displacement and high vibrability significantly exceeds the expectation under random conditions.
Methods
Numerical model
@@ Methods
$U_{a}(r_{ij})=\epsilon\left[a_{0}\left(\frac{\sigma_{ij}}{r_{ij}}\right)^{12}-a_{1}\left(\frac{\sigma_{ij}}{r_{ij}}\right)^{6}+\sum_{l=0}^{3}{c_{2l}\left(\frac{r_{ij}}{\sigma_{ij}}\right)^{2l}}\right].$ (2)
@@
Data availability
The datasets generated and analysed during the current study are available upon reasonable request by contacting the corresponding authors.
Funding
MPC acknowledges support by the Singapore Ministry of Education through grants MOE-T2EP50221-0016 and RG152/23. CJ and MB acknowledge the support of the Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01). MB acknowledges the support of the sponsorship from the Yangyang Development Fund.
Author contributions statement
M.P.C. and M.B. devised the project. M.P.C., S.M. and M.B. wrote the paper. S.M., D.S.Y.H. performed the numerical simulations and prepared the figures. All authors contributed to the scientific discussion and refinement of the manuscript.
Supplementary information
Density of vibrational modes
We compare the density of vibrational modes obtained from Fourier transforming the velocity autocorrelation function (points) and via the direct diagonalisation of the Hessian matrix in Fig. 7 for systems with 256k particles. In the figure, we also illustrate the Debye’s prediction, $D_{\rm Debye}=A_{D}\omega^{D-1}$. In 2D, $A_{D}=\frac{2}{8\pi\rho/(c_{l}^{-2}+c_{s}^{-2})}$, with $\rho$ the number density and $c_{l}=\sqrt{\frac{K+\mu}{\rho}}$ and $c_{s}=\sqrt{\frac{\mu}{\rho}}$ the longitudinal and the transverse shear velocities. In 3D, $A_{D}=\frac{3}{18\pi^{2}\rho/(c_{l}^{-3}+2c_{s}^{-3})}$, $c_{l}=\sqrt{\frac{K+\frac{4\mu}{3}}{\rho}}$ and $c_{s}=\sqrt{\frac{\mu}{\rho}}$. We determine the bulk $K$ and shear $\mu$ moduli by investigating the stress-strain relationships in the linear response regime.
Figure 7: Different approaches to evaluate the vDOS. We compare diverse approaches to evaluating the vDOS, in 2D and 3D. Evaluating the vDOS by Fourier transforming the velocity autocorrelation function and via the diagonalization of the Hessian gives consistent results. Due to our system’s large size, 256k particles, we compute only the lower end of the spectrum via the direct diagonalization of the Hessian.
Cage-relative measures filter plane waves at the boson peak
To evaluate the cage-relative ability to filter extended modes of norm $|{\bm{u}}|^{2}=1$, we consider $|{\bm{u}}^{\rm cr}|^{2}$. Filtering occurs if $|{\bm{u}}^{\rm cr}|^{2}\ll 1$. As extended modes, we consider phonons with various wavelengths. There is no need to consider the superposition of phonons as the cage-relative ‘operator’ is linear, $(a{\bm{u}}+b{\bm{v}})^{\rm cr}=a{\bm{u}}^{\rm cr}+b{\bm{v}}^{\rm cr}$. Any linear combination of ‘filtered’ displacement fields is also filtered.
Fig. 8 illustrates $|{\bm{p}}^{\rm cr}|^{2}$ for phonon ${\bf p}$ of wavelength $\lambda$, in two dimensions. The larger the phonon wavelength, the more efficient the filtering. Importantly, filtering is highly efficient at the wavelength $\lambda_{\rm bp}=\frac{2\pi c_{s}}{\omega_{\rm bp}}$ of the phonons expected to hybridize at the boson peak, as $|{\bm{p}}^{\rm cr}|^{2}$ is in the range $10^{-4}$ to $10^{-3}$, its value depending on the cutoff $x_{c}$ of the interaction potential.
@@
Figure 8: Filtering phonons via the cage-relative measured. Squared norm of the cage-relative displacement field $|{\bm{p}}^{\rm cr}|^{2}$ associated with phonons ($|{\bm{p}}|^{2}=1$) of wavelength $\lambda$, in two dimensions. The vertical lines mark the wavelength $\lambda_{\rm bp}$ of phonons at the boson peak frequency for the extremal investigated $x_{c}$ values. Analogous results hold in 3D.
Density of vibrational defects
We define defects as clusters of particles whose cage-relative displacement is within the top $c=5\%$. We explore how the density of defects, $\rho_{d}$, depends on the percentage $c$ of particles with the highest cage-relative displacement.
As $c\to 0$, $\rho_{d}$ vanishes since there are no clusters, meaning no particles meet the clustering criteria. As $c$ increases, $\rho_{d}$ grows, reflecting both the increasing size and number of clusters. However, beyond a certain point, as clusters begin to merge, $\rho_{d}$ decreases, leading to a maximum in defect density at an intermediate $c$ value.
Fig. 6a shows that, in 2D, the density of defects remains $x_{c}$ independent for $c<10\%$. Unless this threshold is considered unrealistically high, the exact value of $c$ does not significantly affect our conclusions.
In 3D, the behavior of $\rho_{d}$ for small $x_{c}$ values is illustrated in Fig. 6c: $\rho_{d}$ remains unchanged between $x_{c}=1.8$ and $x_{c}=1.7$ and increases as $x_{c}$ decreases further. Thus, our results are not highly sensitive to the exact choice of $x_{c}$.
Figure 9: Cutoff dependence of the density of defects. We perform a cluster analysis of the cof the particles with the largest magnitude of the cage relative displacement (in the mode at boson peak frequency) to identify the defects. a the density of defects $\rho_{d}$ as a maximum as a function of $c$, and is roughly constant as $c$ varies from $5\%$ to $7\%$. b $\rho_{d}$ remains constant in this $c$ range for all $x_{c}$, also when we consider as defects only cluster with at least $n_{\rm min}$ particles. We illustrate the considered values of $n_{\rm min}$ for $x_{c}=1.5$.

### v2

@@ Abstract
Figure 3: 1 Heat map of the magnitude of the coarse-grained cage-relative displacement, normalised by its average value, with superimposed quadrupolar displacements of the defects. The figure refers to $x_{c}=1.7$ and depicts 1/12th of the system’s area. b Density map of the two largest (in modulus) eigenvalues of the deviatoric strain associated with the defects, normalised to its maximum value. The density is higher close to the $\Upsilon_{2}=-\Upsilon_{1}$ line, implying that the quadrupoles correspond to pure shear deformations. c We represent the quadrupolar deformations by attaching to each defect’s centre of mass the displacements along the principal axis $\pm\Upsilon_{1}{\bm{u}}_{1}$ (blue), $\pm\Upsilon_{2}{\bm{u}}_{2}$ (green) and $\pm\Upsilon_{3}{\bm{u}}_{3}$ (red). Since quadrupoles mostly correspond to pure-shear deformations, ${\bf u}_{3}$ is barely visible.
Vibrational properties – To investigate the vibrational properties of these clusters, we first evaluate the strain tensors coarse-grained over a small length scale. Since the strain is insensitive to local uniform translations, as the cage-relative measure, the cage-relative clusters occur in regions of high local strain. This allows us to evaluate a defect strain tensor $\mathcal{E}=\mathcal{E}_{\rm vol}+\mathcal{E}_{\rm dev}$, as detailed in Appendix E. We find that these tensors are dominated by their deviatoric component, indicating that defects perform volume-preserving deformations, as $R=\frac{\sqrt{Q}}{|\rm{Tr}(\mathcal{E})|}\gg 1$. Here, $Q=1/2\sum\Upsilon_{k}^{2}$ is the quadrupolar charge and $\Upsilon_{k}$ are the eigenvalues of the traceless $\mathcal{E}_{\rm dev}$, and $|\rm{Tr}(\mathcal{E})|$ quantifies the degree of volumetric deformations. In the following, we assume $\Upsilon_{1}<0$, order the eigenvalues so that $|\Upsilon_{i}|>|\Upsilon_{i+1}|$, and indicate their associated eigenvectors with ${\bf U}_{k}$. A cluster extends into the directions with $\Upsilon_{k}>0$ while it shrinks in the others for half a period of oscillation, with particles reversing their direction of motion during the remaining half period.
The traceless condition implies $\Upsilon_{1}=-\Upsilon_{2}$ in two dimensions. Figure 3(a) illustrates the quadrupoles associated with each cluster on a heat map of the magnitude of the coarse-grained cage-relative displacement, finding expected correlations. The figure illustrates a small fraction of the system and represents each quadrupole with four arrows with a length proportional to $\pm\Upsilon_{i}{\bm{U}}_{i}$. See Fig. 5 in Appendix A for a snapshot of the whole system and illustrations of the underlying displacement field.
In three dimensions, different deformation modes satisfy the traceless condition $\sum\Upsilon_{i}=0$. We expose the dominant deformation mode by analyzing the probability distribution $P(\Upsilon_{1},\Upsilon_{2})$. In the $\Upsilon_{1}$-$\Upsilon_{2}$ plane, this distribution is supported within a region bounded by the line $\Upsilon_{1}+\Upsilon_{2}=0$, which represents pure shear deformation, and the line $\Upsilon_{1}+2\Upsilon_{2}=0$, which corresponds to volume-preserving isotropic ($\Upsilon_{3}=\Upsilon_{2}=-\Upsilon_{1}/2$) uniaxial compressions. Figure 3(b) shows that quadrupolar deformations are predominantly pure shear, as $P(\Upsilon_{1},\Upsilon_{2})$ peaks near the $\Upsilon_{1}+\Upsilon_{2}=0$ line for different values of $x_{c}$. This result is further supported by direct visualization of the quadrupoles (Figure 3(c)), depicted as collections of 6 arrows with lengths proportional to $\pm\Upsilon_{i}{\bm{U}}_{i}$. Since $|\Upsilon_{3}|\ll|\Upsilon_{2}|$, the blue arrows corresponding to displacements in the $\bm{U}_{3}$} direction are not visible, making the quadrupoles appear as a two-dimensional cross. Overall, these investigations show that modes at the boson peak comprise plane waves hybridized with two-dimensional defects that perform pure shear deformations, in both two- and three-spatial dimensions.
@@
Figure 4: a The coordination number of vibrability clusters increases with $n$, following a trend similar to that of clusters formed based on cage-relative displacement (see Fig.3b). b Particles exhibiting high cage-relative displacement at the boson peak (red) and high vibrability (blue), both in the top 5%, for $x_{c}=1.7$. Strong correlations between the two are observed. c The fraction of particles displaying both large cage-relative displacement and high vibrability significantly exceeds the expectation under random conditions.
The connection with ‘stringlets’ – At this point, one might then wonder how our findings are compatible with the recent claims by Hu and Tanaka [28], who suggested that the BP is induced by one-dimensional clusters, or ‘stringlets’. These are clusters of particles with a large vibrational amplitude, or equivalently, a high ‘vibrability’ $\Psi_{i}=\sum_{k}\frac{1}{\omega_{k}^{2}}\left|{\bf e}_{k,i}\right|^{2}$, where the sum extends over all modes with frequencies $\omega_{k}$ close to the boson peak frequency. We have repeated their analysis in our simulation model. Figure 4(a) shows that the average coordination number of the particles participating in the resulting cluster is large, indicating that these are compact and extended objects, rather than one-dimensional, very similar to the localised defects investigated in Fig. 2(b). In Figure 4(b), we observe a high spatial correlation between particles with vibrability in the top 5% (blue) alongside those with cage-relative displacements in the top 5% (red), indicating a high correlation between vibrational motion and cage-relative displacement at the BP frequency. These correlations are further confirmed by measuring the fraction $f$ of particles with high vibrability and cage-relative displacement, which is much larger than expected in the absence of correlations, as we illustrate in Figure 4(c). Overall, our analysis provides strong evidence that quadrupolar localised defects and high-vibrability clusters are two sides of the same coin. In our systems, however, these clusters are not string-like, and their motion is quadrupolar rather than one-dimensional.
Conclusions – These findings support earlier results [18, 19] suggesting a relation between the boson peak and quasi-localised vibrational modes in a soft-potential model perspective [6, 7, 8, 9]. Our approach has the merit of allowing for a straightforward identification of the hybridized defects in 2D, when these are not visible in the displacement field, and in 3D. The connection between our defects and QLMs deserves further investigation. In particular, while our defects exclusively undergo pure shear deformations, Ref. [49] reported a low-frequency QLM undergoing simple shear deformations where extension occurs in one direction and contraction in the orthogonal ones. In addition, we find that the displacement is predominantly two-dimensional, even in three dimensions, and hence, it does not align with that of the low-frequency localized modes [50].
Along the lines of [18], it would be interesting to understand whether the same defective structures could also be rationalized using shear modulus fluctuations and whether our findings are compatible with another successful theory of the BP – heterogeneous elasticity theory [51]. Moreover, it would be fruitful to investigate the relation between the defects discussed in our manuscript and (i) the geometric charges responsible for plastic screening and anomalous elastic response [52], (ii) the topological vortex-like defects recently connected to plastic soft spots [53, 54].
Acknowledgements- MPC acknowledges support by the Singapore Ministry of Education through grants MOE-T2EP50221-0016 and RG152/23. CJ and MB acknowledge the support of the Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01). MB acknowledges the support of the sponsorship from the Yangyang Development Fund.
Appendix A Extracting defects from a disordered vibrational mode
Fig. 5a presents an eigenmode at the boson peak frequency, characterized by swirls that indicate hybridization with plane waves. This hybridization prevents visually isolating quadrupolar-like defects as recently proposed [19, 20]. Fig. 5b displays the defects identified by our proposed filtering technique, while panel c provides a magnified view of a region covering 1/12th of the system’s area, offering a clearer visualization of the quadrupolar-like localized defects.
Figure 5: a Eigenvector at the boson peak frequency for a two-dimensional system with $x_{c}=1.7$. It is arduous to spot in similar modes isolated quadrupolar-looking defects [19, 20]. b Defects identified by our approach. c A zoom on a region corresponding to 1/12th of the system’s area.
Appendix B Numerical Model
@@
$U_{a}(r_{ij})=\epsilon\left[a_{0}\left(\frac{\sigma_{ij}}{r_{ij}}\right)^{12}\hskip-11.38092pt-a_{1}\left(\frac{\sigma_{ij}}{r_{ij}}\right)^{6}\hskip-5.69046pt+\sum_{l=0}^{3}{c_{2l}\left(\frac{r_{ij}}{\sigma_{ij}}\right)^{2l}}\right].$ (2)
@@
We evaluate the vDOS over a broad frequency range by computing the Fourier transform of the velocity autocorrelation function in the linear response regime by linearizing the equation of motion, implying the absence of anharmonic effects. Additionally, we obtain the vDOS by diagonalizing the Hessian matrix, which provides access to eigenvectors, ${\bf e}_{k}$, and eigenfrequencies, $\omega_{k}$, in the low-frequency regime, including the boson peak. These methods yield consistent results. Fig.6(a) shows the reduced vDOS, $\color[rgb]{0,0,0}D(\omega)/A_{D}\omega^{d-1}$}, with the Debye prediction, $\color[rgb]{0,0,0}D_{\rm Debye}=A_{D}\omega^{d-1}$}, highlighting the BP in two and three spatial dimensions. Here, $A_{D}=d/\omega_{D}^{d}$ where $d$ is dimension, $\omega_{D}^{d}=\frac{18\pi^{2}\rho}{c_{l}^{-3}+2c_{s}^{-3}}$ in 3d and $\omega_{D}^{d}=\frac{8\pi\rho}{c_{l}^{-2}+c_{s}^{-2}}$ in 2d, with the longitudinal and shear velocities given by $c_{l}=\sqrt{\frac{K+\frac{4}{3}\mu}{\rho}}$ in 3d, $c_{l}=\sqrt{\frac{K+\mu}{\rho}}$ in 2d and $c_{s}=\sqrt{\frac{\mu}{\rho}}$ in both 2d and 3d. Here, $K$ is the bulk modulus and $\mu$ is the shear modulus, which we measure by investigating the stress-strain relationships in the linear response regime. In 3D, increasing $x_{c}$ raises the peak amplitude (see Fig. 6b) while lowering its frequency, $\omega_{\rm bp}$ (panel c), suggesting that shorter attractive wells stabilize glasses, as previously proposed [18, 36, 38, 34]. In 2D, the effect of $x_{c}$ is weaker, indicating that the material stability, which depends on $x_{c}$ in our model, has a dimensionality-dependent influence on the BP. It remains to be ascertained whether these results are universal across different systems. For 3D, we analyze $x_{c}\geq 1.5\sigma$, as smaller values maintain a phonon band structure at high frequencies, complicating $\omega_{\rm bp}$ identification, as illustrated in Fig. 7. In 2D, we consider $x_{c}\geq 1.4\sigma$, since for lower values the BP broadens, making $\omega_{\rm bp}$ difficult to estimate.
Figure 6: a The reduced vDOS in three and (inset) two dimensions for systems differing in the extension $x_{c}$ of the intermolecular attractive potential well. In 3D, the materials lose stability as $x_{c}$ increases, as the boson peak intensity increases (b) while its frequency decreases (c). In 2D, attraction does not sensibly influence the excess vibrational modes. Frequencies are expressed in units of the natural frequency $\omega_{0}=c_{s}/a_{0}$, with $a_{0}$ the interparticle spacing and $c_{s}$ the shear wave speed. The inset of (c) illustrates the localised defects that, according to our investigation, induce the boson peak in a three-dimensional system with $x_{c}=1.7$.
Figure 7: Density of states normalized by Debye’s prediction for $x_{c}=1.4$ in 3d. In finite systems, at low frequencies, the discrete nature of the spectrum emerges, and hence, $D(\omega)/A_{D}\omega^{2}$ approaches zero rather than a constant. At small $x_{c}$, the discrete nature of the spectrum extends to higher frequencies, making it difficult to identify the boson peak frequency.
Appendix C Cage-relative filtering approach
To prove that our filtering approach does not filter out localized vibrations, we illustrate in Fig. 8(a) a snapshot of a two-dimensional system where each particle is coloured based on its $|{\bm{e}}_{k,i}|^{2}$, for a low-frequency localised mode $k$. The localised excitation remains visible in panel (b), where particles are coloured according to the magnitude of their cage-relative displacement, $|{\bm{e}}_{k,i}^{\rm cr}|^{2}$, proving that our approach preserves localized defects. Panels (c) and (d) illustrate the same results for a hybridized plane wave.
Figure 8: a Colour map of the magnitude $|{\bm{e}}_{k,i}|^{2}$ of the particle displacement associated with a low-frequency mode with $\omega_{k}=0.23\omega_{\rm bp}$. A localized defect is visible. b The defect remains visible in the colour map of the magnitude of the cage-relative displacement, $|{\bm{e}}_{k,i}^{\rm cr}|^{2}$. In all plots, the magnitudes are scaled so that their average value is 11. c,d report analogous results at a similar frequency for a hybridized rather than a quasi-localized mode.
@@ Appendix C Cage-relative filtering approach
Figure 9: Squared norm of the cage-relative displacement field $|{\bm{p}}^{\rm cr}|^{2}$ associated with phonons ($|{\bm{p}}|^{2}=1$) of wavelength $\lambda$, in two dimensions. The vertical lines mark the wavelength $\lambda_{\rm bp}$ of phonons at the boson peak frequency for the extremal investigated $x_{c}$ values. Analogous results hold in 3D.
To evaluate the cage-relative ability to filter extended modes of norm $|{\bm{u}}|^{2}=1$, we consider $|{\bm{u}}^{\rm cr}|^{2}$. Filtering occurs if $|{\bm{u}}^{\rm cr}|^{2}\ll 1$. As extended modes, we consider phonons with various wavelengths. There is no need to consider the superposition of phonons as the cage-relative ‘operator’ is linear, $(a{\bm{u}}+b{\bm{v}})^{\rm cr}=a{\bm{u}}^{\rm cr}+b{\bm{v}}^{\rm cr}$. Fig. 9 illustrates $|{\bm{p}}^{\rm cr}|^{2}$ for phonon $\bf p$} of wavelength $\lambda$, in two dimensions. Filtering is more efficient at larger wavelengths, where the induced deformation is more affine. Importantly, filtering is highly efficient at the wavelength $\lambda_{\rm bp}=\frac{2\pi c_{s}}{\omega_{\rm bp}}$ of the phonons expected to hybridize at the boson peak, as $|{\bm{p}}^{\rm cr}|^{2}$ is in the range $10^{-4}$ to $10^{-3}$, its value depending on the cutoff $x_{c}$ of the interaction potential.
Appendix D Cluster linear size
Fig. 10 shows that the radius of gyration of the clusters is exponentially distributed, at large $R_{g}$, $P(R_{g})=ae^{-R_{g}/\lambda}$, which allow us to associate a typical length scale $\lambda$ to each system.
Figure 10: Distribution of the defects’ radius of gyration, in 2D and 3D, for diverse values of $x_{c}$.
Appendix E Cluster deviatoric strain
We evaluate the local strain tensor ${\bm{\epsilon}}=\frac{1}{2}\left({\bm{\nabla}}{\bm{e}}+({\bm{\nabla}}\bm{e})^{T}\right),$ associated with the displacement field (eigenvector, $\bm{e}$}) of the mode at the boson peak frequency, coarse-grained on grids of side-length $w\simeq 1.1$. We decompose this tensor into its volumetric and deviatoric components: $\bm{\varepsilon}=\bm{\varepsilon}_{\rm vol}+\bm{\varepsilon}_{\rm dev}$. Here, $\bm{\varepsilon}_{\rm vol}$ represents local volumetric deformations. We define a cluster deviatoric strain as $\mathcal{E}_{\rm dev}=\sum\bm{\varepsilon}_{\rm dev}$, where the sum extends over all grid cells containing particles of the considered cluster.